OpenStudy (anonymous):

Jenna buys 12 lb of meat. Some of the meat is chicken and costs $1.90/lb. The rest is beef, which costs $4.30/lb. She spends a total of $34.80. Which equation models this situation? Let c represent the number of pounds of chicken she buys. A. (1.9 + 4.3)(12 – c) = 34.8 B. 4.3c + 1.9(c + 14) = 34.8 C. 4.3c + 1.9(12) = 34.8 D. 1.9c + 4.3(12 − c) = 34.8

OpenStudy (campbell_st):

well the cost equation is 1.9c + 4.3b = 34.80 the weight equation is c + b = 12 so making b the subject b = 12 - c now substitute this into the cost equation at the top so you can get the correct answer hope it helps

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

wait i dont understand.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so the cost equation of the chicken (c) and beef (b) is 1.9c + 4.3b = 38.40 the weight of the beef is b = 12 - c so substituting this value of b into the cost equation, will all you to have the total cost in terms of chicken(c) 1.9c + 4.3(12 - c) = 34.80 if necessary you could then calculate how much chicken you bought. hope it makes sense